a grocery shopper bought 6 ib of steak and 20 ib of groundbeef the steak cost $5/ib more than the groundbeef and the total bill was $82 what was the price per pound for each type of meat?

Answer by jim_thompson5910(35256)


6 pounds of steak
20 pounds of groundbeef

Steak costs $5 more per pound. If x is the cost per pound of groundbeef, then x+5 is the cost per pound of steak.

(amount of steak)*(cost per pound of steak)+(amount of ground beef)*(cost per pound of ground beef) = total bill
(6 pounds)*(x+5 dollars)+(20 pounds)*(x dollars) = 82
6(x+5)+20x = 82
6x+30+20x = 82
26x+30 = 82
26x+30-30 = 82-30 .......... subtract 30 from both sides
26x = 52
26x/26 = 52/26 .............. divide both sides by 26
x = 2

If x = 2, then,
x+5 = 2+5 = 7

It costs $2 per pound for ground beef, and it costs $7 per pound for steak

Check:
(amount of steak)*(cost per pound of steak)+(amount of ground beef)*(cost per pound of ground beef) = total bill
(6 pounds)*(7 dollars per pound)+(20 pounds)*(2 dollars per pound) = 82
(6*7) + (20*2) = 82
42 + 40 = 82
82 = 82
Solution has been confirmed